OpenStudy (anonymous):

does anyone know how to solve this: x/a = y/b does anyone know how to solve this: x/a = y/b @Mathematics

OpenStudy (anonymous):

what are you solving for here?

OpenStudy (anonymous):

One equation, four variables!!!!! not possible..........

OpenStudy (anonymous):

again, the "x"

OpenStudy (mathteacher1729):

You can express one variable in terms of the other, but you can't "solve" it because that would mean saying something like "x = some number", "y = some number", etc... there is not enough info to solve, is what I'm saying. :)

OpenStudy (anonymous):

x/a = y/b x = ay/b

OpenStudy (anonymous):

x=ay/b

OpenStudy (anonymous):

directions say: "solve each equation for x"

OpenStudy (anonymous):

Hakriat, it is VERY possible :)

OpenStudy (anonymous):

hahaah ^

OpenStudy (anonymous):

Well, it's possible. Anything in math is possible, if not then there is NO solution :)

OpenStudy (anonymous):

LOL! trueee.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@enazukanovic Study harder for your maths exam.........☺

OpenStudy (anonymous):

Thanks MATH teacher!